# Entropic Trade-off Relations in Stochastic Thermodynamics via Replica Markov Processes

### v2

@@ Abstract
Traditional thermodynamic trade-off relations usually apply to quantities that depend linearly on probability distributions. In contrast, many important information-theoretic measures, such as entropies, are nonlinear and thus difficult to analyze with existing frameworks. Motivated by replica ideas in quantum information and spin-glass theory, we introduce replica Markov processes, i.e., Markovian dynamics of $K$ independent, identical copies, and derive trade-off relations that bound relative moments of replica observables in terms of dynamical activity. By choosing appropriate replica observables, these inequalities translate into bounds on nonlinear quantities of the original single process. Focusing on entropic measures of uncertainty, we obtain upper bounds on the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions, and corresponding bounds for diffusion on networks. We also illustrate how replicas can handle other nonlinear functionals such as extreme-value observables. Finally, we extend the construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the bounds are expressed in terms of the quantum dynamical activity. These results provide entropic counterparts to activity-based uncertainty relations and establish a general method for constraining nonlinear information-theoretic quantities in stochastic and quantum thermodynamics.
@@ I Introduction
Over the past decade, thermodynamic trade-off relations have been extensively studied in stochastic and quantum thermodynamics. These relations express a “no free lunch” principle in thermodynamics: faster, more precise operations require greater thermodynamic resources. In the thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13, 14, 15, 16, 17, 18, 19, 20, 21], trajectory observables are considered and lower bounds on their relative variance are derived, where the bounds involve the entropy production and the dynamical activity. Extensions of these uncertainty relations to the quantum domain have been actively studied [17, 18, 21, 22, 23, 24]. Related work has also examined how to compute state-dependent expectation values [25, 26]. What these relations have in common is that the quantities being computed are linear in the underlying probabilities. For example, in thermodynamic uncertainty relations, we are often interested in quantities of the form
@@
where $\Gamma$ denotes a stochastic trajectory, $\mathcal{P}(\Gamma)$ is the trajectory probability, and $f(\Gamma)$ is an observable like a current. Equation (1) is linear with respect to probability $\mathcal{P}(\Gamma)$. Typical examples of Eq. (1) include dissipated heat and stochastic displacement. In contrast, nonlinear functions of the probabilities are generally much harder to compute, although they play a crucial role. A prominent example are generalized entropies such as Rényi and Tsallis, which measure the degree of uncertainty and the spread of information. The thermodynamic uncertainty relation expresses the uncertainty using the variance of an observable. However, in information theory, uncertainty is often quantified using entropy. When the output is categorical rather than numerical, the variance is not a natural measure of uncertainty. Entropy, by contrast, applies directly to categorical outcomes and can be used to quantify uncertainty in this setting. For example, when quantifying the extent that a random walker on a network has propagated over the network, variance is not suitable because the states are labels rather than numerical values. Establishing thermodynamic bounds for these nonlinear functionals remains a challenge because standard frameworks are inherently linked to linear expectations. There is thus a fundamental need for a theoretical bridge that maps nonlinear information measures onto the stochastic thermodynamics.
Inspired by replica methods from quantum information and spin-glass theory, we introduce replica Markovian dynamics comprising $K$ identical copies of the process and derive bounds on quantities that depend nonlinearly on the underlying probability distributions. In quantum information and statistical mechanics, we often need to compute the expectation values of nonlinear physical quantities. For example, purity requires the square of the density operator, $\mathrm{Tr}[\rho^{2}]$. Although purity serves as a key quantity for distinguishing quantum behavior from classical behavior, its experimental measurement is known to be difficult. The swap trick [27] is a technique to measure purity without performing state tomography. The swap trick uses the following identity:
@@
where $\mathrm{SWAP}$ denotes the swap operator. The identity given in Eq. (2) shows that the purity $\mathrm{Tr}[\rho^{2}]$ can be obtained by preparing two replicas of $\rho$ and measuring the expectation value of the SWAP operator acting on $\rho\otimes\rho$. This avoids full state tomography of $\rho$, which would require many measurement shots to reliably estimate $\rho^{2}$. The use of replicas has been highly effective in classical shadow tomography [28, 29, 30] or in the study of the barren plateau in parametrized quantum circuits [31]. The application of hypothetical replicas to the evaluation of nonlinear quantities has a long history. In spin glasses [32, 33, 34], for example, we often need to evaluate the disorder average of the logarithm of the partition function $\mathbb{E}[\ln Z]$, which is difficult to compute directly. The replica method addresses this by introducing $n$ replicas of the system and expressing the logarithm as a limit $\ln Z=\lim_{n\to 0}(Z^{n}-1)/n$. This technique has been widely used in the analysis of learning theory [35, 36, 37, 38, 39] and conformal field theory [40, 41] . These examples illustrate the power of using replicas to calculate expectations of nonlinear quantities.
@@
We introduce a method for deriving trade-off relations for nonlinear quantities of a single Markov process. This is achieved by considering $K$ independent identical copies (replicas) of the process and applying trade-off relations to suitably chosen replica observables. Through this approach, these nonlinear quantities are expressed as linear expectations within the replica framework. Crucially, replicas can be treated as purely virtual, analogous to the replica trick, allowing us to obtain fundamental constraints on the original single process. Our work then focuses on entropic measures of uncertainty. We derive upper bounds on both the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions, relating these to dynamical activity [Eqs. (24) and (25)]. Additionally, we establish Tsallis bounds controlled by the initial activity [Eqs. (27) and (28)] Analogous results also obtained for network diffusion. Specifically, we obtain the time derivative [Eq. (31)] and the value of the Tsallis entropy for state distributions [Eq. (33)], where their bounds comprise the dynamical activity. Furthermore, we establish a bound on the Tsallis entropy of the state distribution, relating it to the initial dynamical activity [Eqs. (34) and (35)]. A notable finding is that these initial activity-controlled bounds depend solely on the local escape rate from the initial node [Eqs. (34) and (35)]. Beyond entropy, we demonstrate the utility of replicas for handling other nonlinear functionals, including extreme-value observables [Eqs. (38)–(41)]. Finally, we generalize this construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the resulting bounds are formulated in terms of quantum dynamical activity [Eq. (54), (55) and (57)]. These results provide entropic counterparts to activity-based uncertainty relations. While conventional uncertainty relations typically lower-bound relative variances by the inverse activity, our bounds similarly show that entropy-based uncertainty is fundamentally constrained by dynamical activity.
@@
Figure 1: Illustration of Markov processes. (a) Single process scenario. This scenario is usually considered in stochastic thermodynamics. For a Markov process, we consider stochastic trajectories, random sample paths that represent individual realizations of the process. We are interested in stochastic quantities associated with the trajectories, whose statistics are linear in the trajectory probability. (b) $K=2$ replica process scenario. We consider two identical Markov processes. The state space is a product space of each Markov process and the trajectories are defined on the product space. We are interested in stochastic quantities of the replica process, which become nonlinear quantities in the original single process.
@@ II.1 Replica Markov processes
where $t_{0}=0$ is the initial time and $X_{0}$ is the initial state. Typically, we consider the counting observable:
$N(\Gamma)=\sum_{\nu,\mu(\nu\neq\mu)}C_{\nu\mu}N_{\nu\mu},$ (5)
where $N_{\nu\mu}$ is the number of jumps from $B_{\mu}$ to $B_{\nu}$ within the interval $0,\tau]$ and $C_{\nu\mu}$ are their associated real weights. As the first-level generalization, we can consider the observable $N_{\circ}(\Gamma)$. $N_{\circ}(\Gamma)$ can be an arbitrary function of $\Gamma$ as long as it satisfies the following condition:
@@
where $\Gamma_{\varnothing}$ represents a trajectory with no jump. Apparently, the counting observable in Eq. (5) satisfies the condition Eq. (6). Let $N_{\bullet}(\Gamma)$ be an arbitrary function defined on $\Gamma$. If we denote the sets of functions $N$, $N_{\circ}$, and $N_{\bullet}$ by $\{N\}$, $\{N_{\circ}\}$, and $\{N_{\bullet}\}$, respectively, then the inclusion holds:
$\{N\}\subset\{N_{\circ}\}\subset\{N_{\bullet}\}.$ (7)
As mentioned in the introduction, we consider replicas of the original process and regard the whole Markov process as a combined process (Fig. 1(b)). For simplicity, here we introduce the case of two replicas. That is, there are two identical processes, where there is no interaction between them. It is straightforward to extend this case to scenarios where there are $K$ replica processes. The state space is a product of the two state sets. The set of states in the combined Markov process of two replicas is given by $\mathfrak{B}^{(\otimes 2)}\equiv\{(B_{1}^{(1)},B_{1}^{(2)}),(B_{1}^{(1)},B_{2}^{(2)}),\ldots,(B_{D}^{(1)},B_{D}^{(2)})\}$. Therefore, the number of total states in the combined process is $|\mathfrak{B}^{(\otimes 2)}|=D^{2}$. From now on, when using a superscript $(k)$, it denotes a quantity associated with the $k$-th process, and when using a superscript $(\otimes K)$, it denotes a quantity for the combined system in which all processes $1,2,\dots,K$ are regarded as a single combined process. When we omit the superscript, the corresponding variables refer to those in the single process (i.e., $K=1$ case).
@@
where $t_{0}=0$ and $X_{0}^{(\otimes 2)}$ is the initial state. For instance, suppose that the trajectories of the first and second processes are
@@ II.2 Replica trade-off relations
Introducing the combined process with two replicas, we will focus on a general scenario in which there are $K$ replicas. We now introduce observables associated with trajectories. In the same way as in the single process case, we can consider observables $N^{(\otimes K)}(\Gamma^{(\otimes K)})$, $N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})$, and $N_{\bullet}^{(\otimes K)}(\Gamma^{(\otimes K)})$. Specifically, $N^{(\otimes K)}$ is the counting observable, which yields the weighted sum of the number of jumps within the interval $0,\tau]$ in the $K$ replica process. The observable $N_{\circ}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})$ satisfies
@@
where $\Gamma_{\varnothing}^{(\otimes K)}$ is a trajectory in the combined process with no jump. The observable $N_{\bullet}^{(\otimes K)}$ is an arbitrary function of $\Gamma^{(\otimes K)}$.
Let $\mathbb{E}[X]$ denote the expectation value of a random variable $X$. Since $\Gamma^{(\otimes K)}$ is a random variable, observable $N^{(\otimes K)}(\Gamma^{(\otimes K)})$ is a random variable as well. From the uncertainty relation in Ref. [5], the following relation holds in the Markov process with $K$ replicas:
$\frac{\mathrm{Var}[N_{\bullet}^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N_{\bullet}^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{A}(\tau)}.$ (14)
@@
Note that the dynamical activity $\mathfrak{a}(t)$ in Eq. (16) is a quantity of a single process [Eq. (3)]; the definition is the same as conventional Markov process settings. Dynamical activity was originally proposed as an order parameter in the study of glass transitions [42]. The dynamical activity plays a central cost term in thermodynamic trade-off relations [3, 43, 5, 44]. Since the replica process comprises $K$ independent processes, the dynamical activity of the combined process is $K\mathcal{A}(\tau)$, which directly yields Eq. (14).
Moreover, by employing another trade-off relation in Ref. [45], we obtain (see Appendix B for the derivation):
@@ II.3 Entropic measures
Figure 2: Illustration of two scenarios considered in trade-off relations, (a) the trajectory observable and (b) the state distribution. (a) Trajectory observable bound in a Markov process. For the Tsallis entropy bound for trajectory observables, we consider a trajectory $\Gamma$ and its associated observables $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. The Tsallis entropy is considered for $N_{\bullet}(\Gamma)$ and $N_{\circ}(\Gamma)$. (b) State distribution bound in the diffusion on a network. A random walker starts from a node in the network, which is represented by dark red. At time $t=\tau$, the random walker diffuses over the network and the probability $p_{\mu}(\tau)$ of the walker position is expressed by red, where darker color corresponds to higher probability. The Tsallis entropy is considered for $p_{\mu}(\tau)$.
By considering specific observables for $N_{\bullet}^{(\otimes K)}$ and $N_{\circ}^{(\otimes K)}$ in Eqs. (14) and (17), we derive trade-off relations for a single process. In particular, we focus on one-parameter generalizations of entropy that quantify different aspects of uncertainty in an arbitrary probability distribution $\{p_{n}\}$. The Rényi entropy [46] of order $\alpha$ $(0<\alpha<\infty)$ is defined as
@@
Varying $\alpha$ continuously interpolates between several standard entropic quantities: in the limit $\alpha\to 0$, $H_{\alpha}^{\mathrm{R}}$ approaches the max-entropy, which depends only on the number of outcomes with nonzero probability; as $\alpha\to 1$, it reduces to the Shannon entropy; for $\alpha=2$, it becomes the collision entropy; and in the limit $\alpha\to\infty$, it converges to the min-entropy determined solely by the largest probability. Another widely used one-parameter generalization is the Tsallis entropy [47, 48], defined by
$H_{q}^{\mathrm{T}}[p_{n}]\equiv\frac{1-\sum_{n}p_{n}^{q}}{q-1},$ (20)
where $q$ is the entropic index. As with Rényi entropy, changing $q$ explores different uncertainty features and connects multiple entropic definitions: in the limit $q\to 0$, $H_{q}^{\mathrm{T}}$ approaches $\mathfrak{W}-1$, where $\mathfrak{W}$ is the number of outcomes with nonzero probability; as $q\to 1$, it reduces to the Shannon entropy; for $q=2$, it yields $H_{2}^{\mathrm{T}}=1-\sum_{n}p_{n}^{2}$, i.e., the complement of the collision probability; and in the limit $q\to\infty$, it tends to 0 for any non-degenerate distribution. For $q>1$, the Tsallis entropy satisfies
$0\leq H_{q}^{\mathrm{T}}<\frac{1}{q-1}\;\;(q>1).$ (21)
Both entropies are related via
@@
The Rényi entropy is widely used in statistics and quantum information theory. In contrast, the Tsallis entropy is mainly applied to non-extensive statistical mechanics and the study of complex systems.

### v3

@@ Abstract
Traditional thermodynamic trade-off relations usually apply to quantities that depend linearly on probability distributions. In contrast, many important information-theoretic measures, such as entropies, are nonlinear and therefore difficult to analyze with existing frameworks. Motivated by replica methods in quantum information and spin-glass theory, we introduce replica Markov processes, i.e., Markovian dynamics of $K$ independent, identical copies, and derive trade-off relations that bound relative moments of replica observables in terms of the dynamical activity. By choosing appropriate replica observables, these inequalities translate into bounds on nonlinear quantities of the original single process. Focusing on entropic measures of uncertainty, we obtain upper bounds on both the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions. Moreover, we derive upper bounds on the Rényi and Tsallis entropies, where the bounds involve the initial dynamical activity. Analogous bounds can be derived for the entropies of the state distribution, which quantify the degree of diffusion. In particular, we provide upper bounds on both the time derivative and the value of the Tsallis entropy of state distributions. Moreover, we show that the Rényi and Tsallis entropies of the state distribution are bounded from above by terms involving the local escape rate from the initial node. We also illustrate how replicas can be used to study other nonlinear functionals such as extreme-value observables. Finally, we extend the construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the bounds are expressed in terms of the quantum dynamical activity. These results provide entropic counterparts to activity-based uncertainty relations and establish a general method for constraining nonlinear information-theoretic quantities in stochastic and quantum thermodynamics.
@@ I Introduction
Over the past decade, thermodynamic trade-off relations have been extensively studied in stochastic and quantum thermodynamics. These relations express a “no free lunch” principle in thermodynamics: faster, more precise operations require greater thermodynamic resources. In the thermodynamic uncertainty relations [1, 2, 3, 4, 5, 6, 7, 8, 9, 10, 11, 12, 13], trajectory observables are considered, and lower bounds on their relative variance are derived, where the bounds involve the entropy production and/or the dynamical activity. Extensions of these uncertainty relations to the quantum domain have been actively studied [14, 15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25, 26, 27]. Related work has also examined how to compute state-dependent expectation values [28, 29]. What these relations have in common is that the quantities being computed are linear in the underlying probabilities. For example, in thermodynamic uncertainty relations, we are often interested in quantities of the form
@@
where $\Gamma$ denotes a stochastic trajectory, $\mathcal{P}(\Gamma)$ is the trajectory probability, and $f(\Gamma)$ is an observable, such as a current. Equation (1) is linear with respect to the probability $\mathcal{P}(\Gamma)$. Typical examples of Eq. (1) include dissipated heat and stochastic displacement. In contrast, nonlinear functions of the probabilities are generally much harder to compute, although they play a crucial role. A prominent example is generalized entropies such as Rényi and Tsallis entropies, which measure the degree of uncertainty and the spread of information. The thermodynamic uncertainty relation expresses uncertainty using the variance of an observable. However, in information theory, uncertainty is often quantified using entropy. When the output is categorical rather than numerical, variance is not a natural measure of uncertainty. Entropy, by contrast, applies directly to categorical outcomes and can be used to quantify uncertainty in this setting. For example, when quantifying the extent to which a random walker on a network has propagated over the network, variance is not suitable because the states are labels rather than numerical values. Establishing thermodynamic bounds for these nonlinear functionals remains a challenge because standard frameworks are inherently tied to linear expectations. There is thus a fundamental need for a theoretical bridge that maps nonlinear information measures onto stochastic thermodynamics.
Inspired by replica methods in quantum information and spin-glass theory, we introduce replica Markov process comprising $K$ identical copies of the process and derive bounds on quantities that depend nonlinearly on the underlying probability distributions. In quantum information and statistical mechanics, we often need to compute the expectation values of nonlinear physical quantities. For example, purity requires the square of the density operator, $\mathrm{Tr}[\rho^{2}]$. Although purity serves as a key quantity for distinguishing quantum behavior from classical behavior, its experimental measurement is known to be difficult. The swap trick [30] is a technique to measure purity without performing state tomography. The swap trick uses the following identity:
@@
where $\mathrm{SWAP}$ denotes the swap operator. The identity given in Eq. (2) shows that the purity $\mathrm{Tr}[\rho^{2}]$ can be obtained by preparing two replicas of $\rho$ and measuring the expectation value of the SWAP operator acting on $\rho\otimes\rho$. This avoids full state tomography of $\rho$, which would require many measurement shots to reliably estimate $\rho^{2}$. The use of replicas has been highly effective in classical shadow tomography [31, 32, 33]. The use of hypothetical replicas for the evaluation of nonlinear quantities has a long history. In spin glasses [34, 35, 36], for example, we often need to evaluate the disorder average of the logarithm of the partition function $\mathbb{E}[\ln Z]$, which is difficult to compute directly. The replica method addresses this by introducing $n$ replicas of the system and expressing the logarithm as the limit $\ln Z=\lim_{n\to 0}(Z^{n}-1)/n$. This technique has been widely used in the analysis of learning theory [37, 38, 39, 40, 41] and conformal field theory [42, 43] . These examples illustrate the power of using replicas to calculate the expectations of nonlinear quantities.
@@
We introduce a method for deriving trade-off relations for nonlinear quantities of a single Markov process. This is achieved by considering $K$ independent, identical copies (replicas) of the process and applying trade-off relations to suitably chosen replica observables. Through this approach, these nonlinear quantities are expressed as linear expectations within the replica framework. Crucially, replicas can be treated as purely virtual, analogous to the replica trick, allowing us to obtain fundamental constraints on the original single process. Our work then focuses on entropic measures of uncertainty. We derive upper bounds on both the time derivative and the value of the Tsallis entropy for general trajectory-observable distributions, relating these to dynamical activity [Eqs. (24) and (25)]. Additionally, we establish upper bounds on the Rényi and Tsallis entropies, which are constrained by the initial dynamical activity [Eqs. (27) and (28)]. Analogous results are also obtained for state distribution of a Markov process. Specifically, we obtain bounds on the time derivative [Eq. (31)] and the value of the Tsallis entropy for state distributions [Eq. (34)], where these bounds are determined by the dynamical activity. Furthermore, we establish bounds on the Rényi and Tsallis entropies of the state distribution, relating them to the initial dynamical activity [Eqs. (35) and (36)]. A notable finding is that these initial-activity-controlled bounds depend solely on the local escape rate from the initial node [Eqs. (35) and (36)]. Beyond entropy, we demonstrate the utility of replicas for handling other nonlinear functionals, including extreme-value observables [Eqs. (40)–(43)]. Finally, we generalize this construction to continuously monitored open quantum systems governed by Lindblad dynamics, where the resulting bounds are formulated in terms of quantum dynamical activity [Eq. (56), (57), and (59)]. These results provide entropic counterparts to activity-based uncertainty relations. While conventional uncertainty relations typically lower-bound relative variances by the inverse activity, our bounds similarly show that entropy-based uncertainty is fundamentally constrained by dynamical activity.
@@
Figure 1: Illustration of Markov processes. (a) Single process scenario with $D=3$ states. This scenario is usually considered in stochastic thermodynamics. For a Markov process, we consider stochastic trajectories, which are random sample paths that represent individual realizations of the process. We are interested in stochastic quantities associated with the trajectories, whose statistics are linear in the trajectory probability. (b) $K=2$ replica process scenario with $D^{2}=9$ states. We consider two identical Markov processes. The state space is the product space of each Markov process, and the trajectories are defined on the product space. We are interested in stochastic quantities of the replica process, which become nonlinear quantities in the original single process.
@@ II.1 Replica Markov processes
where $t_{0}=0$ is the initial time and $X_{0}$ is the initial state. Let $N(\Gamma)$ be an arbitrary real-valued function defined on a trajectory $\Gamma$. We consider a subclass of the observable $N(\Gamma)$, denoted by $N_{\circ}(\Gamma)$. The observable $N_{\circ}(\Gamma)$ satisfies
@@
where $\Gamma_{\varnothing}$ represents a trajectory with no jumps. An important subclass of $N_{\circ}(\Gamma)$ is the counting observable, defined as a linear combination of jump counts:
$N_{\sharp}(\Gamma)\equiv\sum_{\nu,\mu(\nu\neq\mu)}C_{\nu\mu}N_{\nu\mu},$ (6)
where $N_{\nu\mu}$ is the number of jumps from $B_{\mu}$ to $B_{\nu}$ within the interval $0,\tau]$ and $C_{\nu\mu}$ are their associated real weights. Clearly, $N_{\sharp}(\Gamma)$ satisfies Eq. (5). If we denote the sets of functions $N$, $N_{\circ}$, and $N_{\sharp}$ by $\{N\}$, $\{N_{\circ}\}$, and $\{N_{\sharp}\}$, respectively, then the following inclusion holds:
$\{N_{\sharp}\}\subset\{N_{\circ}\}\subset\{N\}.$ (7)
As mentioned in the introduction, we consider replicas of the original process and regard the whole Markov process as a combined process (Fig. 1(b)). For simplicity, here we introduce the case of two replicas. That is, there are two identical processes, with no interaction between them. It is straightforward to extend this case to scenarios where there are $K$ replica processes. The state space is a product of the two state sets. The set of states in the combined Markov process of two replicas is given by $\mathfrak{B}^{(\otimes 2)}\equiv\{(B_{1}^{(1)},B_{1}^{(2)}),(B_{1}^{(1)},B_{2}^{(2)}),\ldots,(B_{D}^{(1)},B_{D}^{(2)})\}$. Therefore, the total number of states in the combined process is $|\mathfrak{B}^{(\otimes 2)}|=D^{2}$. From now on, when using a superscript $(k)$, it denotes a quantity associated with the $k$-th process, and when using a superscript $(\otimes K)$, it denotes a quantity for the combined system in which all processes $1,2,\dots,K$ are regarded as a single combined process. When we omit the superscript, the corresponding variables refer to those in the single process (i.e., $K=1$ case).
@@
where $t_{0}^{(\otimes 2)}=0$ and $X_{0}^{(\otimes 2)}$ is the initial state. For instance, suppose that the trajectories of the first and second processes are
@@ II.2 Replica trade-off relations
Introducing the combined process with two replicas, we will focus on a general scenario in which there are $K$ replicas. We now introduce observables associated with trajectories. In the same way as in the single process case, we can consider observables $N^{(\otimes K)}(\Gamma^{(\otimes K)})$, $N_{\circ}^{(\otimes K)}(\Gamma^{(\otimes K)})$, and $N_{\sharp}^{(\otimes K)}(\Gamma^{(\otimes K)})$. Specifically, The observable $N^{(\otimes K)}$ is an arbitrary function of $\Gamma^{(\otimes K)}$. The observable $N_{\circ}^{(\otimes K)}(\Gamma_{\varnothing}^{(\otimes K)})$ satisfies
@@
where $\Gamma_{\varnothing}^{(\otimes K)}$ is a trajectory in the combined process with no jumps. $N_{\sharp}^{(\otimes K)}$ is the counting observable, which yields the weighted sum of the number of jumps within the interval $0,\tau]$ in the $K$-replica process.
Let $\mathbb{E}[X]$ denote the expectation value of a random variable $X$. Since $\Gamma^{(\otimes K)}$ is a random variable, the observable $N^{(\otimes K)}(\Gamma^{(\otimes K)})$ is also a random variable. From the uncertainty relation in Ref. [5], the following relation holds in the Markov process with $K$ replicas:
$\frac{\mathrm{Var}[N^{(\otimes K)}]}{\tau^{2}\left(\partial_{\tau}\mathbb{E}[N^{(\otimes K)}]\right)^{2}}\geq\frac{1}{K\mathcal{A}(\tau)}.$ (14)
@@
Note that the dynamical activity $\mathfrak{a}(t)$ in Eq. (16) is a quantity of a single process [Eq. (3)]; the definition is the same as that in conventional Markov process settings. Dynamical activity was originally proposed as an order parameter in the study of glass transitions [44]. The dynamical activity plays a central cost term in thermodynamic trade-off relations [3, 45, 5, 46]. Since the replica process comprises $K$ independent processes, the dynamical activity of the combined process is $K\mathcal{A}(\tau)$, which directly yields Eq. (14).
Moreover, by employing another trade-off relation in Ref. [47], we can obtain another relation for the replica process. Recall that $N_{\circ}(\Gamma)$ is the observable that satisfies the condition given in Eq. (5). We obtain (see Appendix C for the derivation):
@@ II.3 Entropic measures
Figure 2: Illustration of two scenarios considered in trade-off relations: (a) the trajectory observable and (b) the state distribution. (a) Trajectory observable bound in a Markov process. For the Tsallis entropy bound for trajectory observables, we consider a trajectory $\Gamma$ and its associated observable $N(\Gamma)$. The entropies are considered for the trajectory observable, and the bounds on the time-derivative and their values are obtained. (b) State distribution bound in the diffusion on a network. In this scenario, a random walker starts from a node in the network, which is represented by dark red. At time $t=\tau$, the random walker diffuses over the network, and the probability $p_{\mu}(\tau)$ of the walker position is expressed by red, where darker color corresponds to higher probability. The entropies are considered for the state distribution $p_{\mu}(\tau)$, and the bounds on the time-derivative and their values are obtained.
By considering specific observables for $N^{(\otimes K)}$ and $N_{\circ}^{(\otimes K)}$ in Eqs. (14) and (17), we derive trade-off relations for a single process. In particular, we focus on one-parameter generalizations of entropy that quantify different aspects of uncertainty in an arbitrary probability distribution $\{p_{n}\}$. One widely used one-parameter generalization is the Tsallis entropy [48, 49], defined by
$H_{q}^{\mathrm{T}}[p_{n}]\equiv\frac{1-\sum_{n}p_{n}^{q}}{q-1},$ (19)
where $q$ is the entropic index. Varying $q$ continuously interpolates between several standard entropic quantities: in the limit $q\to 0$, $H_{q}^{\mathrm{T}}$ approaches $\mathfrak{W}-1$, where $\mathfrak{W}$ is the number of outcomes with nonzero probability; as $q\to 1$, it reduces to the Shannon entropy; for $q=2$, it yields $H_{2}^{\mathrm{T}}=1-\sum_{n}p_{n}^{2}$, i.e., the complement of the collision probability; and in the limit $q\to\infty$, it tends to 0 for any non-degenerate distribution. For $q>1$, the Tsallis entropy satisfies
$0\leq H_{q}^{\mathrm{T}}<\frac{1}{q-1}\;\;(q>1).$ (20)
Another important one-parameter generalization is the Rényi entropy [50] of order $\alpha$ $(0<\alpha<\infty)$, defined as
@@
By varying $\alpha$ continuously, we can interpolate between several standard entropy measures: in the limit $\alpha\to 0$, $H_{\alpha}^{\mathrm{R}}$ approaches the max-entropy, which depends only on the number of outcomes with nonzero probability; as $\alpha\to 1$, it reduces to the Shannon entropy; for $\alpha=2$, it becomes the collision entropy; and in the limit $\alpha\to\infty$, it converges to the min-entropy determined solely by the largest probability. The two entropies are related via
@@
The Tsallis entropy is mainly applied to non-extensive statistical mechanics and the study of complex systems, whereas the Rényi entropy is widely used in statistics and quantum information theory.